# Three short-period Earth-sized planets around M dwarfs discovered by TESS: TOI-5720 b, TOI-6008 b and TOI-6086 b

### v1

@@ 4.3 Ancillary information
Since the three planet candidates have a low SNR, we conducted further analysis other than those provided by DAVE to make our vetting process more solid. TESS has a large pixel scale, about $21^{\prime\prime}$/pixel, with a focus-limited PSF. In order to check whether any resolved or unresolved sources fell within the aperture mask, thus contaminating the photometric signal, we consulted stellar catalogs (Wenger et al. 2000; Gaia Collaboration et al. 2021b). While the centroids module should shed light on the contamination flux due to resolved nearby stars, the same cannot be said for the unresolved sources within the same pixel of the star target. At this level we can only flag the presence of a potentially contaminating source which dilutes the transit depth. The worst scenario is if the contaminating source is the origin of the transit event. We here stress this is an intrinsic limit of this analysis which can be overcome by follow-up photometric campaigns (see Sect. 4.6).
@@
TOI 5720 has only a very faint (TESS magnitude $=18.162$) nearby target, TIC 903650250, which is $\approx 55$ arcsec apart and seven magnitudes dimmer.
@@
On the other hand, TOI 6008 is located in a region of the sky overdense of stars ($82$ known targets in $1$ arcmin). In particular, $5$ stars fall within the same pixel of TOI 6008 even if the brightest one (TIC 286201079) is almost $5$ magnitudes fainter than TOI 6008. Moreover, the brightest star is TIC 286201082 (TESS magnitude $=11.57$) which is $41$ arcsec ($\approx 2$ pixels) away from our target and $\approx 1$ magnitude brighter than it. Thus, the centroid measurements are triggered by the photometric variability of TIC 286201082. Another $42$ arcsec distant star, TIC 286201047 (TESS magnitude $=13$), contributes to make the photocenter measurements challenging. Despite most centroid measurements are unreliable, a few are not altered by nearby sources. For those, the centroid is aligned with the TOI 6008’s position.
@@
The photometry obtained from TOI 6086 observations are contaminated by the star light of TIC 18318284 (TESS magnitude $=13.25$) which is only one magnitude fainter than TOI 6086 and $46$ arcsec distant. Moreover, the photocenter measurements are completely counterfeit due to the presence of a $120^{\prime\prime}$ arcsec star (Gaia DR3 4597447931261454080) which is 2 magnitudes brighter than TOI 6086. Also in this case, a few clean centroid measurements do not show any flag for an offset.
@@
We also decided to conduct a Pixel Level light curve (PLL) analysis using the publicly available LATTE [^4] LATTE: https://github.com/noraeisner/LATTE pipeline developed by Eisner (2022). This test allowed us to inspect the light curve for each pixel in the field of view of the corresponding target pixel file to check whether the transit occurs in nearby pixels of the mask. None of the pixels within the aperture mask other than that corresponding to TOI 6008 and TOI 6086 shows a reliable dip in the starlight. We repeated the same procedure for all the transits observed by TESS with the same conclusions.
@@
Despite the orbital path of TESS has been designed to maximize the sky coverage while reducing the number of obstructions, the TESS Full Frame Images can occasionally be contaminated primarily by the zodiacal light and scattered light from solar system objects (Gangestad et al. 2013; Sullivan et al. 2015). Thus, during each TESS sector’s observation, the flux of the background can vary widely. If any unusual events in the background occur at the transit time of the planet candidate, they could introduce stilted signals into the light curve, contaminating or, even worse, originating the detected transit. In order to ensure that no anomaly contaminated the signal, we visually inspected a 3-day-long section of the background flux centered at the time of transit. We found no evidences of background artifacts for TOI 5720.01, TOI 6008.01 and TOI 6086.01.
@@
We found no known TESS momentum dumps in correspondence of the transits of TOI 5720.01, TOI 6008.01 and TOI 6086.01, thus ruling out any possible instrumental systematic scenario.
@@ 4.4 Archival imaging
We used archival images of TOI-5720, TOI-6008 and TOI-6086 to exclude the possibility of any background unresolved stellar companion that could be blended with targets at these current positions. Such objects might be introducing the transits signals that we detected in our observations, and skew the results that we obtained from the global analysis.
TOI-5720 has a high proper motion of 350.4 $mas/yr$. We obtained images from POSS-II/DSS (Minkowski & Abell 1963) on 1995, POSS-II/DSS (Lasker et al. 1996) on 1999 and PanSTARRS-1 (Chambers et al. 2016) on 2011 in the blue, infrared and $z_{s}$ bands, respectively, and spanning 28 years with our current observations. TOI-6008 has a relative low proper motion of 158.4 $mas/yr$. We obtained images from POSS-II/DSS (on 1989 in blue band), POSS-II/DSS (on 1994 in the infrared band) and PanSTARRS-1 (on 2011 in the $z_{s}$ band), and spanning 34 years with our current observations. TOI-6086 has also a relative low proper motion of 110.6 $mas/yr$. We also obtained images from POSS-I/DSS (on 1952 in red band), POSS-II/DSS (on 1995 in the infrared band) and PanSTARRS-1 (on 2011 in the $z_{s}$ band), and spanning 71 years with our current observations. TOI-5720, TOI-6008 and TOI-6086 have moved only by $\sim 9.8$, $\sim 5.4$ and $\sim 7.8$ arcsec between 1995 and 2023, 1989 and 2023, and 1952 and 2023, respectively. There is no source in the current day positions of TOI-5720, TOI-6008 and TOI-6086 in any of these archival images (see Figure 9).
Refer to captionRefer to captionRefer to caption
Figure 9: Archive images of TOI-5720 (top panel), TOI-6008 (middle panel) and TOI-6086 (bottom panel). The previous and current positions of the targets are shows in red and cyan colors, respectively. Top row (TOI-5720): 1995 blue image from POSS-II/DSS, 1999 infrared image from POSS-II/DSS2, 2011 $z_{s}$ image from PanSTARRS1, and 2023 $z_{s}$ image from MuSCAT3 (shown from left to right). Middle row (TOI-6008): 1989 blue image from POSS-II/DSS, 1994 infrared image from POSS-II/DSS2, 2011 $z_{s}$ image from PanSTARRS1, and 2023 $z_{s}$ image from MuSCAT3 (shown from left to right). Bottom row (TOI-6086): 1952 red image from POSS-I/DSS, 1995 infrared image from POSS-II/DSS2, 2011 $z_{s}$ image from PanSTARRS1, and 2023 $z_{s}$ image from MuSCAT3 (shown from left to right).
@@ 4.5 Statistical validation
We also employed the TRICERATOPS pipeline (Giacalone et al. 2021) to assess the likelihood of false positive scenarios. TRICERATOPS calculates the probability that a light curve originates from a transiting planet or a wide range of other scenarios in a Bayesian framework. In particular, it returns the False Positive Probability (FPP), which is the probability that the observed transit is not caused by a planet (e.g., an eclipsing binary) and the Nearby False Positive Probability (NFPP), which is the probability that the signal comes from a nearby resolved star. Moreover, TRICERATOPS also allows to implement, if available, the high-contrast imaging observations in order to better constrain the calculation. Typically, a planet candidate is considered statistically validated if it returns a FPP$<0.01$ and a NFPP $<0.001$ (Morton 2012, 2015; Giacalone et al. 2021). We ran TRICERATOPS using the light curves of TOI 5720.01, TOI 6008.01 and TOI 6086.01 phase-folded on the orbital periods obtained by the photometric modelling, as we will show in Sect.6. We constraint the overall calculation by implementing the high-contrast imaging observations mentioned in Sect. 4.4. Using ground-based observations (see Sect.4.6), we detected the source of a transit signal on-target, and thus we excluded other nearby sources, which are obviously not contaminating the transit signal, from our FPP calculations, and that is the reason we obtained NFPP $=0$ for the three candidates. We obtained FPP $=(9.6\pm 0.4)\times 10^{-3}$ for the signal associated with TOI 5720.01, FPP $=(0.2\pm 0.05)\times 10^{-3}$ for TOI 6008.01 and FPP $=(5.00\pm 0.05)\times 10^{-3}$ for TOI 6086.01.
@@ 4.6 Ground-based follow-up photometric validation
The ground-based photometric follow-up was carried out on the basis of two objectives: i) identify the source of the transit event; ii) chromaticity check for the transit depth in different wavelengths. The closest neighboring stars to TOI-5720 is TIC 903650250 at $54.7\arcsec$ with a $T_{\rm mag}$ of 18.16, and a $\Delta T_{\rm mag}$ of 6.56. The closest neighboring stars to TOI-6008 are respectively TIC 286201110 at $7.3\arcsec$ with a $T_{\rm mag}$ of 17.32, and a $\Delta T_{\rm mag}$ of 5.02, TIC 286201086 at $12.4\arcsec$ with a $T_{\rm mag}$ of 17.43, and a $\Delta T_{\rm mag}$ of 5.13, TIC 286201079 at $14.5\arcsec$ with a $T_{\rm mag}$ of 16.98, and a $\Delta T_{\rm mag}$ of 4.68, and TIC 286201077 at $16.7\arcsec$ with a $T_{\rm mag}$ of 17.59, and a $\Delta T_{\rm mag}$ of 5.29. The closest neighboring stars to TOI-6086 are respectively TIC 18318294 at $25.3\arcsec$ with a $T_{\rm mag}$ of 16.23, and a $\Delta T_{\rm mag}$ of 3.83, and TIC 18318295 at $33.0\arcsec$ with a $T_{\rm mag}$ of 16.81, and a $\Delta T_{\rm mag}$ of 4.41. See the TESS FOV in Figure 1.
@@
We used uncontaminated photometric aperture sizes of only a few arc-seconds using ground-based telescopes, in order to confirm the transit events on the expected target stars for TOI-5720 b, TOI-6008 b and TOI-6086 b. We collected photometric observations for all targets in different bands (zs, Sloan-i’, Sloan-r’, Sloan-g’), spanning from 477 to 1100 nm. We measured a matching transit depth within $1\sigma$ in all bands. The transit depths obtained in different bands for TOI-5720 b, TOI-6008 b, and TOI-6086 b are presented in Figure 10 and Table 4.
@@
Figure 10: Measured transit depths in different filters (colored dots with error bars) obtained in the global analysis for TOI-5720 b (top panel), TOI-6008 b (middle panel) and TOI-6086 b (bottom panel). The dashed line corresponds to the depth obtained from the achromatic fit in the global analysis with $1\sigma$ error bar (shaded region). All measurements agree with the common transit depth at $1\sigma$.
5 Planet searches using the TESS photometry
@@
Using the TESS data described in Section 2.1 and our SHERLOCK package (Pozuelos et al. 2020; Demory et al. 2020) we aimed to (i) recover the candidates TOI-5720.01, TOI-6008.01, and TOI-6086.01 and (ii) search for extra transiting planets in these systems that may remain unnoticed due to detection thresholds of the SPOC. The SHERLOCK pipeline is specially set up for searching low-S/N transit-like features that might be attributable to planets. This pipeline is used to explore space-based data produced by the TESS and Kepler/K2 missions; for example, it is used in the SPECULOOS project (Sebastian et al. 2021), where is searching for transiting planets orbiting ultra-cool dwarfs and the FATE project (Van Grootel et al. 2021), where is searching for planetary remnants orbiting hot subdwarfs (Thuillier et al. 2022). In particular, it uses a multi-detrend approach applying a bi-weight filter several times with different window sizes and the TLS algorithm (Hippke & Heller 2019) to optimize the transit search by converging to the signals with the highest S/N and SDE (signal-detection-efficiency) (see, e.g., Delrez et al. 2022; Pozuelos et al. 2023). In all the cases, we first found the signal corresponding to the planetary candidate alerted by SPOC; that is, we successfully recovered all the TOIs studied in this work. In subsequent runs, for TOI-5720 and TOI-6086, we found no other signal that might be attributable to planets. On the other hand, for TOI-6008, we found a weak signal corresponding to an Earth-sized planet of $\sim$1.12 R with an orbital period of 16.35 days. The semiautomatic vetting conducted by SHERLOCK did not find any hint of systematics that might be producing the signal; however, the statistical validation (automatically carried out by SHERLOCK but relaying in TRICERATOPS) yielded an NFPP and FPP of 0.175 and 0.356, placing it as a likely false positive (Giacalone et al. 2021). Still, we decided to conduct ground-based observations to fully confirm this signal’s false positive character. We conducted three observations with MuSCAT2 on August 03 2023 and one with MuSCAT3 on August 19 2023; both of the cases, we do not find any evidence of a transit feature in the data, allowing us to conclude that this signal is indeed a false positive. Moreover, if the planet was present, we would see a clear transit in the MuSCAT3 light curve (See Figure 11) with a signal-to-noise ratio of $S/N\sim 6$. Finally, we computed also the Lomb–Scargle periodogram (Lomb 1976; Scargle 1982), there were no indications of flaring activity or stellar modulation seen in the TESS data of these three systems. This implies that the rotational periods of stars are probably longer than the TESS observation window for a single sector.
@@
Figure 11: LCO-HAL-2m0/MuSCAT photometry of TOI-6008.22 collected on UT August 20, 2023in the Sloan-$i^{\prime}$ filter with an exposure time of 30s. The blue data points show the raw light curve. The red data points show the light curve after the transit injection of an Earth-sized planet of $\sim 1.12$ R with the best-fitting model is superimposed in black. The green region shows the predicted ingress and egress of the transit. The light curves are shifted along the y-axis for clarity.
6 Global modelling: Photometric and RVs

### v2

@@ 4.3 Ancillary information
Since the three planet candidates have a low S/N, we conducted analysis in addition to those provided by DAVE to make our vetting process more solid. TESS has a large pixel scale, about $21^{\prime\prime}$/pixel, with a focus-limited PSF. In order to check whether any resolved or unresolved sources fell within the aperture mask, thus contaminating the photometric signal, we consulted stellar catalogs (Wenger et al. 2000; Gaia Collaboration et al. 2021b). While the centroids module should shed light on the contamination flux due to resolved nearby stars, the same cannot be said for the unresolved sources within the same pixel of the star target. At this level, we can only flag the presence of a potentially contaminating source that dilutes the transit depth. The worst scenario is when the contaminating source is the origin of the transit event. We here stress that this is an intrinsic limit of this analysis that can be overcome by follow-up photometric campaigns (see Sect. 4.6).
@@
TOI-5720 has only a very faint (TESS magnitude $=18.162$) nearby target, TIC 903650250, which is $\sim 55$ arcsec apart and seven magnitudes dimmer.
@@
On the other hand, TOI-6008 is located in a region of the sky that is overdense in stars ($82$ known targets in $1$ arcmin). In particular, $5$ stars fall within the same pixel of TOI-6008 even if the brightest star (TIC 286201079) is almost $5$ magnitudes fainter than TOI-6008. Moreover, the brightest star is TIC 286201082 (TESS magnitude $=11.57$), which is $41$ arcsec ($\sim 2$ pixels) away from our target and $\sim 1$ magnitude brighter than it. Thus, the centroid measurements are triggered by the photometric variability of TIC 286201082. Another star at a distance of $42$ arcsec, TIC 286201047 (TESS magnitude $=13$), contributes to make the photocenter measurements challenging. Even though most centroid measurements are unreliable, a few are not altered by nearby sources. For these, the centroid is aligned with the position of TOI-6008.
@@
The photometry obtained from TOI-6086 observations is contaminated by the star light of TIC 18318284 (TESS magnitude $=13.25$), which is only one magnitude fainter than TOI-6086 and lies at a distance of $46$ arcsec. Moreover, the photocenter measurements are completely compromised due to the presence of a $120$ arcsec star (Gaia DR3 4597447931261454080) that is 2 magnitudes brighter than TOI-6086. In this case as well, a few clean centroid measurements do not show any flag for an offset.
@@
We also decided to conduct a pixel level light-curve (PLL) analysis using the publicly available LATTE [^4] LATTE: https://github.com/noraeisner/LATTE pipeline developed by Eisner (2022). This test allowed us to inspect the light curve for each pixel in the field of view of the corresponding target pixel file to check whether the transit occurs in nearby pixels of the mask. None of the pixels within the aperture mask other than the pixel corresponding to TOI-6008 and TOI-6086 shows a reliable dip in the starlight. We repeated the same procedure for all the transits observed by TESS with the same conclusions.
@@
Although the orbital path of TESS has been designed to maximize the sky coverage while reducing the number of obstructions, the TESS full frame images can occasionally be contaminated primarily by the zodiacal light and scattered light from Solar System objects (Gangestad et al. 2013; Sullivan et al. 2015). Thus, during the observation of each TESS sector, the flux of the background can vary widely. If any unusual events in the background occurred at the transit time of the planet candidate, they could introduce stilted signals into the light curve, contaminating or, even worse, triggering the detected transit. In order to ensure that no anomaly contaminated the signal, we visually inspected a three-day-long section of the background flux centered at the time of transit. We found no evidence of background artifacts for TOI-5720.01, TOI-6008.01, and TOI-6086.01.
@@
We found no known TESS momentum dumps in correspondence of the transits of TOI-5720.01, TOI-6008.01, and TOI-6086.01. We therefore rule out any possible instrumental systematic scenario.
@@ 4.4 Archival imaging
We used archival images of TOI-5720, TOI-6008, and TOI-6086 to exclude any background-unresolved stellar companion that could be blended with targets at these current positions. Objects like this might introduce the transits signals that we detected in our observations, and skew the results that we obtained from the global analysis.
TOI-5720 has a high proper motion of 350.04 $mas/yr$. We obtained images from POSS-II/DSS (Minkowski & Abell 1963) in 1995, POSS-II/DSS (Lasker et al. 1996) in 1999, and PanSTARRS-1 (Chambers et al. 2016) in 2011 in the blue, infrared,and $z_{s}$ bands, respectively, and spanning 28 years with our current observations. TOI-6008 has a relatively low proper motion of 158.3 $mas/yr$. We obtained images from POSS-II/DSS (in 1989 in blue band), POSS-II/DSS (in 1994 in the infrared band), and PanSTARRS-1 (in 2011 in the $z_{s}$ band), and spanning 34 years with our current observations. TOI-6086 also has a relatively low proper motion of 110.6 $mas/yr$. We also obtained images from POSS-I/DSS (in 1952 in red band), POSS-II/DSS (in 1995 in the infrared band), and PanSTARRS-1 (in 2011 in the $z_{s}$ band), and spanning 71 years with our current observations. TOI-5720, TOI-6008, and TOI-6086 have only moved by $\sim 9.8$, $\sim 5.4$, and $\sim 7.8$ arcsec between 1995 and 2023, 1989 and 2023, and 1952 and 2023, respectively. There is no source in the current day positions of TOI-5720, TOI-6008, and TOI-6086 in any of these archival images (see Figure 9).
Figure 9: Archive images of TOI-5720 (top panel), TOI-6008 (middle panel), and TOI-6086 (bottom panel). The previous and current positions of the targets are shown in red and cyan, respectively. Top row (TOI-5720): 1995 blue image from POSS-II/DSS, 1999 infrared image from POSS-II/DSS2, 2011 $z_{s}$ image from PanSTARRS1, and 2023 $z_{s}$ image from MuSCAT3 (shown from left to right). Middle row (TOI-6008): 1989 blue image from POSS-II/DSS, 1994 infrared image from POSS-II/DSS2, 2011 $z_{s}$ image from PanSTARRS1, and 2023 $z_{s}$ image from MuSCAT3 (shown from left to right). Bottom row (TOI-6086): 1952 red image from POSS-I/DSS, 1995 infrared image from POSS-II/DSS2, 2011 $z_{s}$ image from PanSTARRS1, and 2023 $z_{s}$ image from MuSCAT3 (shown from left to right).
@@ 4.5 Statistical validation
We also employed the TRICERATOPS pipeline (Giacalone et al. 2021) to assess the likelihood of false-positive scenarios. TRICERATOPS calculates the probability that a light curve originates from a transiting planet or from a wide range of other scenarios in a Bayesian framework. In particular, it returns the false-positive probability (FPP), which is the probability that the observed transit is not caused by a planet (e.g., an eclipsing binary), and the nearby false-positive probability (NFPP), which is the probability that the signal comes from a nearby resolved star. Moreover, TRICERATOPS also allows us to implement, if available, the high-contrast imaging observations in order to better constrain the calculation. Typically, a planet candidate is considered statistically validated when it returns an FPP$<0.01$ and a NFPP $<0.001$ (Morton 2012, 2015; Giacalone et al. 2021). We ran TRICERATOPS using the light curves of TOI-5720.01, TOI-6008.01, and TOI-6086.01 phase-folded on the orbital periods obtained by the photometric modelling, as we will show in Sect.6. We constrained the overall calculation by implementing the high-contrast imaging observations mentioned in Sect 4.4. Using ground-based observations (see Sect.4.6), we detected the source of a transit signal on-target, and thus we excluded other nearby sources, which clearly do not contaminate the transit signal, from our FPP calculations. We therefore obtained NFPP $=0$ for the three candidates. We obtained FPP $=(9.6\pm 0.4)\times 10^{-3}$ for the signal associated with TOI-5720.01, FPP $=(0.2\pm 0.05)\times 10^{-3}$ for TOI-6008.01, and FPP $=(5.00\pm 0.05)\times 10^{-3}$ for TOI-6086.01.
@@ 4.6 Ground-based follow-up photometric validation
The ground-based photometric follow-up was carried out on the basis of two objectives: i) identify the source of the transit event, and ii) chromaticity check for the transit depth in different wavelengths. The closest neighboring star to TOI-5720 is TIC 903650250 at $54.7\arcsec$ with a $T_{\rm mag}$ of 18.16, and a $\Delta T_{\rm mag}$ of 6.56. The closest neighboring stars to TOI-6008 are TIC 286201110 at $7.3\arcsec$ with a $T_{\rm mag}$ of 17.32 and a $\Delta T_{\rm mag}$ of 5.02, TIC 286201086 at $12.4\arcsec$ with a $T_{\rm mag}$ of 17.43 and a $\Delta T_{\rm mag}$ of 5.13, TIC 286201079 at $14.5\arcsec$ with a $T_{\rm mag}$ of 16.98, and a $\Delta T_{\rm mag}$ of 4.68, and TIC 286201077 at $16.7\arcsec$ with a $T_{\rm mag}$ of 17.59 and a $\Delta T_{\rm mag}$ of 5.29. The closest neighboring stars to TOI-6086 are TIC 18318294 at $25.3\arcsec$ with a $T_{\rm mag}$ of 16.23, and a $\Delta T_{\rm mag}$ of 3.83, and TIC 18318295 at $33.0\arcsec$ with a $T_{\rm mag}$ of 16.81 and a $\Delta T_{\rm mag}$ of 4.41 (see the TESS FOV in Figure 1).
@@
We used uncontaminated photometric aperture sizes of only a few arcseconds using ground-based telescopes in order to confirm the transit events on the expected target stars for TOI-5720 b, TOI-6008 b and TOI-6086 b. We collected photometric observations for all targets in different bands (zs, Sloan-i’, Sloan-r’, and Sloan-g’), spanning from 477 to 1100 nm. We measured a matching transit depth within $1\sigma$ in all bands. The transit depths obtained in different bands for TOI-5720 b, TOI-6008 b, and TOI-6086 b are presented in Figure 10 and Table 4.
@@
Figure 10: Measured transit depths in different filters (colored dots with error bars) obtained in the global analysis for TOI-5720 b (top panel), TOI-6008 b (middle panel), and TOI-6086 b (bottom panel). The dashed line corresponds to the depth obtained from the achromatic fit in the global analysis with a $1\sigma$ error bar (shaded region). All measurements agree with the common transit depth at $1\sigma$.
5 Planet searches using TESS photometry
@@
Using the TESS data described in Section 2.1 and our SHERLOCK package (Pozuelos et al. 2020; Demory et al. 2020), we aimed to (i) recover the candidates TOI-5720.01, TOI-6008.01, and TOI-6086.01 and (ii) search for additional transiting planets in these systems that may remain unnoticed due to detection thresholds of the SPOC. The SHERLOCK pipeline is especially set up for searching low-S/N transit-like features that might be attributable to planets. This pipeline is used to explore space-based data produced by the TESS and Kepler/K2 missions; for example, it is used in the SPECULOOS project (Sebastian et al. 2021), where it searches for transiting planets orbiting ultracool dwarfs, and the FATE project (Van Grootel et al. 2021), where it searches for planetary remnants orbiting hot subdwarfs (Thuillier et al. 2022). In particular, it uses a multidetrend approach applying a biweight filter several times with different window sizes and the TLS algorithm (Hippke & Heller 2019) to optimize the transit search by converging to the signals with the highest S/N and signal-detection-efficiency (SDE) (see, e.g., Delrez et al. 2022; Pozuelos et al. 2023). In all the cases, we first found the signal corresponding to the planetary candidate to which SPOC alerted; that is, we successfully recovered all the TOIs studied in this work. In subsequent runs, for TOI-5720 and TOI-6086, we found no other signal that might be attributable to planets. On the other hand, for TOI-6008, we found a weak signal corresponding to an Earth-sized planet of $\sim$1.12 R with an orbital period of 16.35 days. The semiautomatic vetting conducted by SHERLOCK did not find any indication of systematics that might produce the signal; however, the statistical validation (automatically carried out by SHERLOCK but relaying in TRICERATOPS) yielded an NFPP and FPP of 0.175 and 0.356, which means that it is a likely false positive (Giacalone et al. 2021). We still decided to conduct ground-based observations to fully confirm the false-positive character of this signal. We conducted three observations with MuSCAT2 on August 3 2023 and one observation with MuSCAT3 on August 19 2023. We found no evidence of a transit feature in the data, and therefore conclude that this signal is indeed a false positive. Moreover, if the planet were present, we would see a clear transit in the MuSCAT3 light curve (See Figure 11) with a signal-to-noise ratio of $S/N\sim 6$. Finally, we also computed the Lomb–Scargle periodogram (Lomb 1976; Scargle 1982), which showed no indications of flaring activity or stellar modulation in the TESS data of these three systems. This implies that the rotational periods of the stars are probably longer than the TESS observation window for a single sector.
@@
Figure 11: LCO-HAL-2m0/MuSCAT photometry of TOI-6008.22 collected on UT August 20, 2023 in the Sloan-$i^{\prime}$ filter with an exposure time of 30s. The blue data points show the raw light curve. The red data points show the light curve after the transit injection of an Earth-sized planet of $\sim 1.12$ R with the best-fitting model superimposed in black. The green region shows the predicted ingress and egress of the transit. The light curves are shifted along the y-axis for clarity.
6 Global modeling: Photometry and radial velocities